OpenStudy (anonymous):

**MEDAL** Solve. -7x-3y=-25 -6x-4y=-20

OpenStudy (anonymous):

\[-6(7x)-3y=-25 and -7(-6x)4y=-20\]

OpenStudy (anonymous):

you - the two equation so you wil just have y.

OpenStudy (anonymous):

Okay, so is −6(7x)−3y=−25and−7(−6x)4y=−20 the answer? This was the last step up a question I didn't understand :l I got the beginning though!

OpenStudy (j-cross18):

Yes.

OpenStudy (anonymous):

Thanks, guys :P

OpenStudy (anonymous):

no haha:) its like (-42x -3y= -25) + (42 x -4y=20) each term is together so -7y=-5

OpenStudy (anonymous):

Oh, i see now!

OpenStudy (anonymous):

so you can find your y and after, you put this y in one of the two equations and you find the x :)

OpenStudy (anonymous):

Thanks! :D